# Centro de Física do Porto e Departamento de Física e Astronomia, Faculdade de Ciências da Universidade do Porto, Rua do Campo Alegre 687, 4169-007, Porto, Portugal τCPHT, CNRS, École Polytechnique, Institut Polytechnique de Paris, Palaiseau, France

### v1

@@ Appendix A RG instability of $SU(2)_{3}\times SU(2)_{3}/SU(2)_{6}$ coset CFT
In this appendix we will show that the $SU(2)_{3}\times SU(2)_{3}/SU(2)_{6}$ coset, which has categorical symmetry $\mathcal{C}$, has a relevant singlet operator under $\mathcal{C}$. First, we need identify how the symmetries in this model act. The complex conjugation $\mathbb{Z}_{2}$ acts on primary operators through
@@
The remaining symmetries are generated by Verlinde lines $\mathcal{L}_{(\ell_{3},\ell_{3}^{\prime},\ell_{6})}$ which are in one-to-one correspondence with coset primaries. The diagonal spin-flip $\mathbb{Z}_{2}$ line $\eta$ is generated by
@@
and are exchanged by the action of the complex conjugation $\mathbb{Z}_{2}$. The action of such Verlinde line on coset primaries is well-known to be given by Verlinde:1988sn
@@
and is therefore a relevant singlet, from which we conclude that $SU(2)_{3}\times SU(2)_{3}/SU(2)_{6}$ is not the stable fixed point of our model, but can potentially be reached by tuning an additional $\mathcal{C}$ invariant coupling, i.e. it is a candidate to be the CFT in our phase diagram.

### v3

@@ A.1 RG instability of $SU(2)_{3}\times SU(2)_{3}/SU(2)_{6}$ coset CFT
We will now show that the $SU(2)_{3}\times SU(2)_{3}/SU(2)_{6}$ coset, which has categorical symmetry $\mathcal{C}$, has a relevant singlet operator under $\mathcal{C}$. First, we need to identify how the symmetries in this model act. The complex conjugation $\mathbb{Z}_{2}$ acts on primary operators through
@@
The remaining symmetries are generated by Verlinde lines $\mathcal{L}_{(\ell_{3},\ell_{3}^{\prime},\ell_{6})}$ which are in one-to-one correspondence with coset primaries $\phi_{(\ell_{3},\ell_{3}^{\prime},\ell_{6})}$. The diagonal spin-flip $\mathbb{Z}_{2}$ line $\eta$ is generated by
@@
and are exchanged by the action of the complex conjugation $\mathbb{Z}_{2}$. The action of such Verlinde line on coset primaries is well-known to be given by Verlinde (1988)
@@
and is therefore a relevant singlet, from which we conclude that $SU(2)_{3}\times SU(2)_{3}/SU(2)_{6}$ is not the stable fixed point of our model, but can potentially be reached by tuning an additional $\mathcal{C}$ invariant coupling, i.e. it is the obvious candidate to be the CFT2 in our phase diagram.
Appendix B Finite size correction to the effective central charge
In this appendix, we discuss how to improve estimates to the effective central charge by performing a finite size scaling analysis. To estimate these finite size corrections we consider a toy beta function
$l\frac{d}{dl}g=-\omega g+\mathcal{O}(g^{2}).$ (74)
Such a beta function naturally appears in perturbation theory with a weakly relevant operator. Perturbative RG flow is a gradient flow, and one can define a potential Jack and Osborn (1990)
$A(g)=-\frac{1}{2}\omega g^{2}+\mathcal{O}(g^{2}).$ (75)
This allows us to write $\beta(g)=\frac{dA(g)}{dg}$, as expected from a gradient flow. Solving the beta function, we get that
$A(t)=2\omega a\cdot e^{-2\omega t}+a_{\infty},\quad{\rm with}\quad t=\log(l).$ (76)
Similar to the central charge in 2D CFT, such a potential AA is also a measure of the effective degrees of freedom, and in fact is proportional to it in two dimensions Zamolodchikov (1986); Cardy (1988). It is therefore reasonable to assume that the central charge has the same finite size corrections. The effective central charge is the derivative of the entanglement entropy with respect to $t=\log(l)$. We therefore have
$EE(t)=a\cdot e^{-2\omega t}+c_{\infty}t+b,$ (77)
where $a,b$ are non-universal constant, while $c_{\infty}$ and $\omega$ depend on the CFT data. The constant $c_{\infty}$ is just the central charge, while $\omega$ is controlled by the leading irrelevant operator $\epsilon^{\prime}$ (invariant under the Fibonacci symmetry), that is $\omega=D-\Delta_{\epsilon^{\prime}}$. We can test this using a single anyon chain with $J>0$. The result in Fig. 24 confirms that the entanglement entropy indeed follows the proposed finite-size behavior.
(a) Entanglement entropy
(b) Effective central charge
Figure 24: Left: The Entanglement entropy of a single anyon chain, fitted using $EE(t)=a\cdot e^{-2t}+c_{\infty}t+b$, which corresponds to the red line. The effective central charge on the right plot is defined as $c_{\rm eff}(t)=\frac{dEE}{dt}$. Right: The effective central charge fitted using $c(t)=c_{\infty}+ae^{-2\omega t}$. The data points are obtained by doing a linear fit of eight neighboring points in the left plot and extracting the slope. The three points near the boundary are discarded because they are likely to be affected by the UV details of the lattice. We obtain $c_{\infty}=0.69998(4)$ and $\omega=1.01(4)$.